$SPAD/src/input s17aef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
showArrayValues true
showScalarValues true
result:=s17aef(0.5,-1)

———-

References^{[1] nothing}

4